The question is :Is the following work correct? In complete sentences, explain why or why not.

3x+4=6x-2

Step 1: x+4=2x-2

Step 2: -x=-6

Step 3: x=6

Found 2 solutions by math_tutor2020, josgarithmetic:
Answer by math_tutor2020(3817)


Step 1 is not correct. You cannot divide solely the 3x by 3 without doing the same thing to the 4 on the left hand side. The same idea applies to the right hand side as well.

If you're going to divide something by 3, then you have to do the same to both sides. In other words, you have to divide each term by 3.

This is one way to do the steps correctly
3x+4 = 6x-2
3x-6x = -2-4
-3x = -6
x = -6/(-3)
x = 2

Check:
3x+4 = 6x-2
3*2+4 = 6*2-2
6+4 = 12-2
10 = 10
We get the same thing on both sides to fully verify the solution.

Answer by josgarithmetic(39617)

------------------------
3x+4=6x-2
Step 1: x+4=2x-2
Step 2: -x=-6
Step 3: x=6
-----------------------

Step #1 is wrong. Also if step #1 was chosen to be a multiplication (using multiplicative inverse), it was done so incorrectly.

A much better choice for step #1 would be ADD negative 3x to both sides...